Find aba-1 where (i) a = (5, 7, 9) , b = (1, 2, 3) (ii) a = (1, 2,5)(3, 4) , b = (1, 4, 5) .


1
Expert's answer
2022-01-24T14:24:16-0500

Let us find "aba^{-1}" where

(i) "a=(5,7,9),\ b=(1,2,3)"


It follows that "a^{-1}=(5,9,7)." Taking into account that the circles "a" and "b" are independent, and hence they commute, we conclude that

"aba^{-1}=(5,7,9) (1,2,3)(5,9,7)=(1,2,3)(5,7,9)(5,9,7)=(1,2,3)=b."



(ii) "a=(1,2,5)(3,4),\ b =(1,4,5)"


It follows that "a^{-1}=(3,4)(1,5,2)." Therefore,

"aba^{-1}=((1,2,5)(3,4))(1,4,5)((3,4)(1,5,2))=((134)(2,5))((3,4)(1,5,2))\n=(123)."